# SelfAI: Building a Self-Training AI System with LLM Agents

### v1

@@ Appendix B Details of SelfAI Benchmark
Figure 6: The relationship between different parameter settings and accuracy. a A hexbin plot showing the joint density of learning rate (log-scaled) and accuracy. b Box plots illustrating the distribution of dropout rates used for models of different depths. c Mean accuracy and standard error across logarithmically binned batch sizes. d Performance comparison of different architectural configurations (layers and units). e A scatter plot of weight decay against momentum, colored by accuracy. f The optimization trajectory, showing accuracy improvement over successive trials, with a moving average trend line. g A correlation matrix quantifying linear relationships between all parameters and the accuracy.
@@
B.1 Failure Cases
Comparison of Best Result Hit Rate. Table 3 compares the experimental results of traditional solvers and LLM-based solvers. Early-stopping LLM-ES partially alleviates redundant trials but remains less stable and less efficient than SelfAI-integrated variants. In contrast, SelfAI consistently enhances trajectory-level reasoning across all LLM families, reducing unnecessary exploration (lower $\text{AUP}_{D}$), improving stopping behavior (lower Stop-Time), and yielding higher overall Score. Mid-sized models such as Qwen2.5-7B and DeepSeek-R1-32B demonstrate particularly favorable exploration and exploitation trade-offs. Although LLMs possess advanced reasoning capabilities and can identify optimal results more efficiently, we found that LLM-based solvers still face challenges related to context limitations and computational fragility.
1) Context limitations. The LCBench dataset exceeds the maximum token limit of GPT-family models, which is why SelfAI does not employ GPT4-o3-mini or GPT4-o3. In practice, search spaces rarely contain thousands of trials, and this issue can be mitigated in future work. For instance, memory management and Retrieval-Augmented Generation (RAG) can help reduce token requirements by introducing additional steps to the experimental recommendation workflow, such as embedding, chunking, and retrieval.
2) Computational fragility. Table 3 compares the best result hit rates across solvers. The performance of LLM-based solvers is highly sensitive to the reasoning strategy, and early stopping often leads to a decline in hit rate. In the supplementary tables, we report the best result $t_{\text{best}}$ for each solver. $t_{\text{best}}=1$ indicates that the optimal result was not found. Moreover, when $t_{\text{best}}=1$, the score metric improves significantly, highlighting the solver’s weakness in optimal stopping. As illustrated in Fig. 11, computational fragility also hinders reliable performance improvement through parameter tuning, as seen in the non-monotonic performance of the DeepSeek-R1 model series. Table 6 summarizes the metrics corresponding to Fig. 11.
Fig. 11 provides a detailed visualization of solver behavior. The Qwen2.5 models consistently identify high-performing configurations early in the search, with Qwen2.5-7B reaching near-optimal regions within the first 10–15 trials. Their trajectories exhibit stable refinement and limited oscillation, suggesting consistent integration of prior experimental evidence. In contrast, the DeepSeek-R1 models show pronounced non-monotonic behavior. DeepSeek-R1-7B frequently oscillates between high and low values, reflecting sensitivity to minor variations in its reasoning process. DeepSeek-R1-70B achieves early improvements but then terminates prematurely, failing to further refine promising regions. Although the GS and BS reach high-performance regions, their trajectories lack the rapid breakthroughs observed in LLM-based solvers. Overall, these trajectory comparisons provide visual evidence of differences in exploration stability, breakthrough efficiency, and stopping behavior across solver families, complementing the aggregate metrics reported in the main text.
@@ Appendix C Supplementary Tables
GS 0.1100 1.000000 0.780000 1.000000 16.63 13
BS 0.329545 0.761200 0.477273 0.863636 16.63 11
Qwen2.5-7b 0.602273 0.487619 0.363636 0.431818 16.63 2
Qwen2.5-14b 0.454545 0.708497 0.272727 0.818182 16.63 4
Qwen2.5-32b 0.374388 0.538240 0.613636 0.636364 16.60 9
Qwen2.5-72b 0.408757 0.536178 0.636364 0.545455 16.60 5
DeepSeek-r1-7b 0.693182 0.318656 0.227273 0.386364 16.63 1
DeepSeek-r1-14b 0.408534 0.471791 0.704545 0.477273 16.60 6
DeepSeek-r1-32b 0.312477 0.198417 1.000000 0.340909 15.77 12
DeepSeek-r1-70b 0.374555 0.576330 0.590909 0.659091 16.60 8
Llama3.3-70b 0.380344 0.285865 0.590909 0.295455 10.79 7
GPT4-o3-mini 0.477273 0.517709 0.454545 0.590909 16.63 3
GPT4-o3 0.363636 0.671366 0.477273 0.795455 16.63 10
@@
Table 9: Performance comparison of different solvers for Mask AutoEncoders (MAE) [69].
@@
Table 11: Performance comparison of different solvers for the image classification task on the LCBench dataset reported from [71].
@@
Table 12: Performance comparison of different solvers for the medical image segmentation task on the BTCV dataset reported from the nnUnet benchmark [73].
@@
Table 13: Performance comparison of different solvers for the nnUnet model on the BraTS dataset [82].
@@
Table 15: Performance comparison of different solvers for bioactivity prediction on Chagas EP20 dataset [95]

### v2

@@ B.5 Drug Discovery
Figure 5: The relationship between different parameter settings and accuracy. a, A hexbin plot showing the joint density of learning rate (log-scaled) and accuracy. b, Box plots illustrating the distribution of dropout rates used for models of different depths. c, Mean accuracy and standard error across logarithmically binned batch sizes. d, Performance comparison of different architectural configurations (layers and units). e, A scatter plot of weight decay against momentum, colored by accuracy. f, The optimization trajectory, showing accuracy improvement over successive trials, with a moving average trend line. g A correlation matrix quantifying linear relationships between all parameters and the accuracy.
@@
B.6 Details of limitations
1) Context limitations. LCBench contains approximately 2000 trials, constituting a large and highly sparse search space. Supplementary Fig. 5 provides a structural analysis of this landscape. Performance depends on interacting hyperparameters such as weight decay, hidden-layer width, and depth, and high-performing regions are sharply localized (Supplementary Fig. 5a-e). In practice, such large and sparse search spaces are common in scientific discovery tasks and can exceed the maximum context length supported by current GPT-family models.
2) Computational fragility. Supplementary Table 3 compares the best result hit rates across solvers. The performance of LLM-based solvers is highly sensitive to the reasoning strategy, and early stopping often leads to a decline in hit rate. In the supplementary tables, we report the best result $t_{\text{best}}$ for each solver. $t_{\text{best}}=1$ indicates that the optimal result was not found. Moreover, when $t_{\text{best}}=1$, the score metric improves significantly, highlighting the solver’s weakness in optimal stopping. As illustrated in Fig. 10, computational fragility also hinders reliable performance improvement through parameter tuning, as seen in the non-monotonic performance of the DeepSeek-R1 model series.
3) Insufficient trajectory integration. Supplementary Fig. 10 provides representative trajectory visualizations that highlight differences in evidence integration and stopping behavior. The Qwen2.5 models typically identify high-performing configurations early, with Qwen2.5-7B reaching near-optimal regions within the first 10-15 trials, followed by stable refinement with limited oscillation. In contrast, DeepSeek-R1 trajectories are more frequently non-monotonic. DeepSeek-R1-7B often oscillates between high- and low-performing regions, indicating instability in how accumulated evidence is leveraged, while DeepSeek-R1-70B can terminate prematurely after early improvements, leaving promising regions under-refined. Classical baselines (GS and BS) may also reach high-performing regions, but their improvements tend to be slower and less punctuated by rapid breakthroughs. Overall, these examples complement the aggregate metrics in the main text by illustrating how limitations can manifest as reduced trajectory stability, delayed breakthroughs, or suboptimal stopping decisions.
@@ Appendix C Supplementary Tables
GS 0.1100 1.0000 0.7800 1.0000 16.63 13
BS 0.3295 0.7612 0.4773 0.8636 16.63 11
Qwen2.5-7b 0.6023 0.4876 0.3636 0.4318 16.63 2
Qwen2.5-14b 0.4545 0.7085 0.2727 0.8182 16.63 4
Qwen2.5-32b 0.3743 0.5382 0.6136 0.6364 16.60 9
Qwen2.5-72b 0.4088 0.5362 0.6364 0.5455 16.60 5
DeepSeek-r1-7b 0.6932 0.3187 0.2273 0.3864 16.63 1
DeepSeek-r1-14b 0.4085 0.4718 0.7045 0.4773 16.60 6
DeepSeek-r1-32b 0.3125 0.1984 1.0000 0.3409 15.77 12
DeepSeek-r1-70b 0.3746 0.5763 0.5909 0.6591 16.60 8
Llama3.3-70b 0.3803 0.2859 0.5909 0.2955 10.79 7
GPT4-o3-mini 0.4773 0.5177 0.4545 0.5909 16.63 3
GPT4-o3 0.3636 0.6714 0.4773 0.7955 16.63 10
@@
Table 9: Performance comparison of different solvers for Mask AutoEncoders (MAE) [mae].
@@
Table 11: Performance comparison of different solvers for the image classification task on the LCBench dataset reported from [zimmer2021auto].
@@
Table 12: Performance comparison of different solvers for the medical image segmentation task on the BTCV dataset reported from the nnUnet benchmark [nnunet_revisit].
@@
Table 13: Performance comparison of different solvers for the nnUnet model on the BraTS dataset [brats].
@@
Table 15: Performance comparison of different solvers for bioactivity prediction on Chagas EP20 dataset [Chagas]